Jane plans to retire in 30 years and would like to have $500,000 in her account. If the interest on her account is 3.92% compounded monthly, how much should she put into her account every month? A. 1,633.33 B. 1,388.89 C. 730.74 D. 653.33 Which would result in the need for a larger monthly deposit to a retirement account with the same ending balance? A. Half the number of years before the money is withdrawn from the account B. Half the interest rate
1
Expert's answer
2021-10-25T15:19:26-0400

"A=\frac{P[(1+\frac{r}{n})^{nt}-1]}{\frac{r}{n}}"

A"=" $ 500,000

P"=?"

r"=" "\frac{3.92}{1200}" Compounded monthly

n"=" 12

t"=" 30 years


"500,000=\frac{P[(1+\frac{3.92}{1200})^{12×30}-1]}{\frac{3.92}{1200}}"

"500,000=684.24P"

"P=730.74"

Half of the years we will use 15 years:

"500,000=\frac{P[(1+\frac{3.92}{1200})^{12×15}-1]}{\frac{3.92}{1200}}"

"500,000=244.49P"

"P=2045.09"

For half the interest we use 1.96% so:

"500,000=\frac{P[(1+\frac{1.96}{1200})^{12×30}-1]}{\frac{1.96}{1200}}"

"500,000=489.50P"

P=1021.45

Half the number of years will result in a larger monthly deposit of 2045.09